OpenStudy (anonymous):

what is f' for \[9^{x ^{4}}\]

OpenStudy (anonymous):

\[(\ln9)4x^39^{x^4}\]

OpenStudy (anonymous):

could it be 4(ln 9)x^3 9^x^3?

OpenStudy (anonymous):

Why not 9^x^4?

OpenStudy (anonymous):

wolf says that the answer is d/dx(9^(x^4)) = 4×9^(x^4) x^3 log(9) but im trying to work out why

OpenStudy (anonymous):

because I don't see that as a possible choice :(

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

a is exactly what I've written

OpenStudy (anonymous):

oh, it's the equivlent to the other one.

OpenStudy (anonymous):

ok its the derivative of the outside times the derivative of the inside: inside one is easy: d/dx x^4 is 4x^3 times the derivative of the outside function d/dx 9^x which is .... (9^x)ln 9....ok I see

OpenStudy (anonymous):

Let \[f=9^{x^4} \], then \[\ln(f)=x^4\ln(9)\], use implicit differentiation.